# RealCamo: Boosting Real Camouflage Synthesis with Layout Controls and Textual-Visual Guidance

### v2

@@ 2 Related Work
Camouflaged Image Generation. Camouflaged image generation, introduced by Chu et al. (2010), can be categorized into two paradigms: Style transfer-based background-guided approaches modify foreground appearance to blend targets into given backgrounds. DCI (Zhang et al., 2020) employs an attention-aware camouflage loss to suppress salient cues while preserving subtle perceptual features, and LCGNet (Li et al., 2022b) achieves extreme camouflage by fusing high-level foreground and background representations. In contrast, out-painting-based foreground-guided methods generate backgrounds conditioned on the foreground target. LAKE-RED (Zhao et al., 2024) introduces a knowledge retrieval–augmented framework using a pretrained codebook, while FACIG (Chen et al., 2025) refines feature integration to reduce foreground distortion. CamoAny (Das and Gopalakrishnan, 2025) further improves background synthesis through controlled out-painting and representation engineering. Despite these advances, existing methods still struggle to simultaneously achieve strong visual similarity and semantic consistency. More recently, the text-driven method CT-CIG (Qian et al., 2025) has shown promise in realistic camouflage synthesis, but the inability to provide aligned annotations limits its applicability to downstream camouflaged object detection applications.
@@ 3.5 Textual-Visual Condition Generator
Figure 4: Qualitative comparison with existing SOTA methods including style transfer-based (column 3 to 4), out-painting-based (column 5 to 9), and text-driven (column 10) approaches. Our results are demonstrated in the rightmost column with red box.
@@
LAKE-RED $_{\textit{CVPR 2024}}$ OP 39.55 0.0212 0.2208 0.9191 88.70 0.0428 0.2881 1.4268 102.67 0.0555 0.2736 1.4216 64.27 0.0355 0.2608 1.2558
FACIG $_{\textit{ICME 2025}}$ OP 27.61 0.0099 82.23 0.0326 96.94 0.0503 52.87 0.0229
CamoAny $_{\textit{CVPR 2025}}$ OP 32.05 0.0140 0.1591 0.2948 52.24 0.0256 0.2052 0.5905 86.35 0.0430 0.1953 0.6173 42.85 0.0233 0.1865 0.5009
CT-CIG $_{\textit{AAAI 2026}}$ TD 30.59 0.0085 0.1972 1.0199 81.60 0.0230 0.1910 1.1938 104.46 0.0241 0.2091 1.4424 52.88 0.0169 0.1991 1.2187
RealCamo (Ours) OP 9.06 0.0015 0.3927 0.4648 11.16 0.0028 0.4670 0.9238 16.11 0.0056 0.4285 0.8365 6.93 0.0025 0.4294 0.7417
@@ 3.7 Background-Foreground Distribution Divergence
$\times$ $\times$ 46.52 0.0149 0.1446 0.7371 73.38 0.0266 0.2002 1.5530 102.35 0.0340 0.1953 1.3155 46.21 0.0212 0.1800 1.2019
$\surd$ $\times$ 37.59 0.0121 0.2910 0.4019 43.97 0.0151 0.3525 0.9349 61.92 0.0280 0.3160 0.8637 27.71 0.0162 0.3198 0.7335
$\times$ $\surd$ 32.98 0.0099 0.1712 0.7405 55.01 0.0137 0.2320 1.5997 87.27 0.0237 0.2186 1.3464 33.14 0.0106 0.2073 1.2289
$\surd$ $\surd$ 31.77 0.0071 0.3254 0.3398 35.61 0.0083 0.4027 0.8659 44.90 0.0143 0.3695 0.7938 20.30 0.0083 0.3659 0.6665
@@
$\surd$ $\times$ $\times$ 32.99 0.0118 0.1917 0.6343
$\surd$ $\surd$ $\times$ 25.30 0.0104 0.2887 0.6994
$\surd$ $\times$ $\surd$ 25.54 0.0102 0.2949 0.6957
$\surd$ $\surd$ $\surd$ 20.30 0.0083 0.3659 0.6665
@@ 4.1 Experimental Settings
Simple $\times$ 46.21 0.0212 0.1800 1.2019
Detailed $\times$ 27.71 0.0162 0.3198 0.7335
$\times$ $\surd$ 33.14 0.0106 0.2073 1.2289
Detailed $\surd$ 20.30 0.0083 0.3659 0.6665
@@ 4.2 Comparison with the State-of-the-arts (SOTAs)
We compare our RealCamo with 12 SOTAs, including five style transfer-based methods (AB (Martino et al., 2016), CI (Chu et al., 2010), AdaIN (Huang and Belongie, 2017), DCI (Zhang et al., 2020), and LCGNet (Li et al., 2022b)), six out-painting-based methods (TFill (Zheng et al., 2022), LDM (Rombach et al., 2022), RePaint-L (Lugmayr et al., 2022), LAKE-RED (Zhao et al., 2024), FACIG (Chen et al., 2025), and CamoAny (Das and Gopalakrishnan, 2025)), and one recently text-driven method (CT-CIG (Qian et al., 2025)). We re-evaluate the benchmark results provided by LAKE-RED and CT-CIG. Due to a lack of available resources, we reproduce CamoAny with the same settings as ours, and directly copy the results from the original paper for FACIG.
@@
Quantitative Comparison. Quantitative results are summarized in Tab. 1. Compared to early style transfer-based methods, recent out-painting-based and text-driven approaches achieve obviously better image quality, indicated by lower FID and KID. Owing to the introduction of explicit layout controls and textual–visual multimodal conditioning, our RealCamo attains highly competitive generation performance. Although it does not achieve the lowest $\mathrm{KL}_{BF}$ score, our method balances visual camouflage with structural and semantic fidelity, producing more realistic images. By revisiting Fig. 4, methods with lower $\mathrm{KL}_{BF}$ values (e.g., CamoAny and TFill) indeed yield stronger visual camouflage, further validating our proposed $\mathrm{KL}_{BF}$ as an effective metric for quantifying camouflage effectiveness.
User Study. We further conduct a user study to comprehensively evaluate the quality of synthesized camouflaged images based on human feedback. Detailed experimental settings, evaluation criteria, and statistical results and analyses are reported in Appendix E.
@@ C.2 Ablations on Textual-Visual Condition
$\surd$ $\times$ $\times$ 35.29 0.0099 0.1629 0.3191 55.49 0.0146 0.2116 0.8249 87.31 0.0251 0.2007 0.7589 32.99 0.0118 0.1917 0.6343
$\surd$ $\surd$ $\times$ 32.94 0.0085 0.2627 0.3565 43.50 0.0116 0.3299 0.9049 55.91 0.0221 0.2735 0.8367 25.30 0.0104 0.2887 0.6994
$\surd$ $\times$ $\surd$ 33.76 0.0086 0.2421 0.3459 43.60 0.0127 0.3421 0.8957 56.74 0.0229 0.3006 0.8454 25.54 0.0102 0.2949 0.6957
$\surd$ $\surd$ $\surd$ 31.77 0.0071 0.3254 0.3398 35.61 0.0083 0.4027 0.8659 44.90 0.0143 0.3695 0.7938 20.30 0.0083 0.3659 0.6665
@@
Simple $\times$ 37.59 0.0121 0.2910 0.4019 43.97 0.0151 0.3525 0.9349 61.92 0.0280 0.3160 0.8637 27.71 0.0162 0.3198 0.7335
Detail $\times$ 37.37 0.0116 0.2951 0.3639 42.94 0.0123 0.3610 0.9039 56.94 0.0235 0.3297 0.8209 26.74 0.0148 0.3286 0.6962
$\times$ $\surd$ 33.66 0.0080 0.3077 0.3591 40.92 0.0109 0.3443 0.8862 50.47 0.0191 0.3483 0.8142 23.44 0.0105 0.3334 0.6865
Detail $\surd$ 31.77 0.0071 0.3254 0.3398 35.61 0.0083 0.4027 0.8659 44.90 0.0143 0.3695 0.7938 20.30 0.0083 0.3659 0.6665

### v3

@@ 2 Related Work
Camouflaged Image Generation. Camouflaged image generation, introduced by Chu et al. (2010), can be categorized into two paradigms: Style transfer-based background-guided approaches modify foreground appearance to blend targets into given backgrounds. DCI (Zhang et al., 2020) employs an attention-aware camouflage loss to suppress salient cues while preserving subtle perceptual features, and LCGNet (Li et al., 2022b) achieves extreme camouflage by fusing high-level foreground and background representations. In contrast, out-painting-based foreground-guided methods generate backgrounds conditioned on the foreground target. LAKE-RED (Zhao et al., 2024) introduces a knowledge retrieval–augmented framework using a pretrained codebook, while FACIG (Chen et al., 2025) reduces foreground distortion with a foreground-aware denoising loss. CamoAny (Das and Gopalakrishnan, 2025) further improves background synthesis through controlled out-painting and representation engineering. Despite these advances, existing methods still struggle to simultaneously achieve strong visual similarity and semantic consistency. Recent text-driven method CT-CIG (Qian et al., 2025) has shown promise in realistic camouflage synthesis, but the inability to provide aligned annotations limits its applicability to downstream camouflaged object detection applications.
@@ 3.5 Textual-Visual Condition Generator
Figure 4: Qualitative comparison with existing SOTA methods including style transfer-based (column 3 to 4), out-painting-based (column 5 to 10), and text-driven (column 11) approaches. Our results are demonstrated in the rightmost column with red box.
@@
LAKE-RED $_{\textit{CVPR 2024}}$ OP 39.55 0.0212 0.2208 0.9191 89.08 0.0435 0.2881 1.4268 102.45 0.0624 0.2736 1.4216 64.33 0.0357 0.2608 1.2558
FACIG $_{\textit{ICME 2025}}$ OP 27.61 0.0099 0.2443 0.8953 82.23 0.0326 0.3234 1.5152 96.96 0.0504 0.2983 1.2428 52.87 0.0229 0.2887 1.2177
CamoAny $_{\textit{CVPR 2025}}$ OP 34.41 0.0129 0.1591 0.2948 79.38 0.0314 0.2052 0.5905 99.06 0.0539 0.1953 0.6173 55.40 0.0248 0.1865 0.5009
CT-CIG $_{\textit{AAAI 2026}}$ TD 30.59 0.0085 0.1972 1.0199 81.60 0.0231 0.1910 1.1938 104.46 0.0241 0.2091 1.4424 52.88 0.0153 0.1991 1.2187
RealCamo (Ours) OP 14.56 0.0021 0.3927 0.4648 91.26 0.0276 0.4670 0.9238 125.65 0.0475 0.4285 0.8365 52.00 0.0173 0.4294 0.7417
@@ 3.7 Background-Foreground Distribution Divergence
$\times$ $\times$ 39.69 0.0144 0.1446 0.7371 103.16 0.0372 0.2002 1.5530 141.34 0.0798 0.1953 1.3155 70.81 0.0317 0.1800 1.2019
$\surd$ $\times$ 36.75 0.0130 0.2910 0.4019 101.28 0.0364 0.3525 0.9349 139.78 0.0738 0.3160 0.8637 65.58 0.0290 0.3198 0.7335
$\times$ $\surd$ 35.42 0.0108 0.1712 0.7405 100.34 0.0346 0.2320 1.5997 138.15 0.0660 0.2186 1.3464 63.97 0.0268 0.2073 1.2289
$\surd$ $\surd$ 31.63 0.0075 0.3254 0.3398 98.28 0.0324 0.4027 0.8659 136.95 0.0613 0.3695 0.7938 60.40 0.0243 0.3659 0.6665
@@
$\surd$ $\times$ $\times$ 63.73 0.0272 0.1917 0.6343
$\surd$ $\surd$ $\times$ 61.53 0.0258 0.2887 0.6994
$\surd$ $\times$ $\surd$ 61.25 0.0250 0.2949 0.6957
$\surd$ $\surd$ $\surd$ 60.40 0.0243 0.3659 0.6665
@@ 4.1 Experimental Settings
Simple $\times$ 65.58 0.0290 0.3198 0.7335
Detailed $\times$ 63.09 0.0277 0.3286 0.6962
$\times$ $\surd$ 61.66 0.0260 0.3334 0.6865
Detailed $\surd$ 60.40 0.0243 0.3659 0.6665
@@ 4.2 Comparison with the State-of-the-arts (SOTAs)
We compare our RealCamo with 12 SOTAs, including five style transfer-based methods (AB (Martino et al., 2016), CI (Chu et al., 2010), AdaIN (Huang and Belongie, 2017), DCI (Zhang et al., 2020), and LCGNet (Li et al., 2022b)), six out-painting-based methods (TFill (Zheng et al., 2022), LDM (Rombach et al., 2022), RePaint-L (Lugmayr et al., 2022), LAKE-RED (Zhao et al., 2024), FACIG (Chen et al., 2025), and CamoAny (Das and Gopalakrishnan, 2025)), and one recently text-driven method (CT-CIG (Qian et al., 2025)). We re-evaluate the benchmark results provided by LAKE-RED, FACIG, and CT-CIG. Due to a lack of available resources, we reproduce CamoAny with the same settings as ours.
@@
Quantitative Comparison. Quantitative results are summarized in Tab. 1. Compared to early style transfer-based methods, recent out-painting-based and text-driven approaches achieve obviously better image quality, indicated by lower FID and KID. Owing to the introduction of explicit layout controls and textual–visual multimodal conditioning, our RealCamo attains highly competitive generation performance. Although it does not achieve the lowest $\mathrm{KL}_{BF}$ score, our method balances visual camouflage with structural and semantic fidelity, producing more realistic images. By revisiting Fig. 4, methods with lower $\mathrm{KL}_{BF}$ values (e.g., CamoAny and TFill) indeed yield stronger visual camouflage, further validating our proposed $\mathrm{KL}_{BF}$ as an effective metric for quantifying camouflage effectiveness. We also note that RealCamo attains higher FID under salient & general object transfer scenes. This is largely due to the domain gap introduced by structural controls. However, this does not indicate degraded perceptual quality, while KID, extensive visualizations, and user study results consistently show that RealCamo produces high-fidelity camouflaged images. These observations highlight the limitations of FID and KID for evaluating camouflaged image generation, which further motivates our introduction of SSIM and $\mathrm{KL}_{BF}$ metrics.
User Study. We further conduct a user study to comprehensively evaluate the quality of synthesized camouflaged images based on human feedback. Detailed evaluation settings, results, and analyses are reported in Appendix E.
@@ C.2 Ablations on Textual-Visual Condition
$\surd$ $\times$ $\times$ 35.39 0.0103 0.1629 0.3191 101.54 0.0355 0.2116 0.8249 137.89 0.0685 0.2007 0.7589 63.73 0.0272 0.1917 0.6343
$\surd$ $\surd$ $\times$ 33.14 0.0085 0.2627 0.3565 100.36 0.0339 0.3299 0.9049 137.16 0.0665 0.2735 0.8367 61.53 0.0258 0.2887 0.6994
$\surd$ $\times$ $\surd$ 32.84 0.0082 0.2421 0.3459 99.01 0.0336 0.3421 0.8957 136.54 0.0620 0.3006 0.8454 61.25 0.0250 0.2949 0.6957
$\surd$ $\surd$ $\surd$ 31.63 0.0075 0.3254 0.3398 98.28 0.0324 0.4027 0.8659 136.95 0.0613 0.3695 0.7938 60.40 0.0243 0.3659 0.6665
@@
Simple $\times$ 36.75 0.0130 0.2910 0.4019 100.28 0.0364 0.3525 0.9349 139.78 0.0738 0.3160 0.8637 65.58 0.0290 0.3198 0.7335
Detail $\times$ 34.63 0.0111 0.2951 0.3639 100.78 0.0353 0.3610 0.9039 138.84 0.0701 0.3297 0.8209 63.09 0.0277 0.3286 0.6962
$\times$ $\surd$ 31.68 0.0081 0.3077 0.3591 99.62 0.0333 0.3443 0.8862 136.82 0.0621 0.3483 0.8142 61.66 0.0260 0.3334 0.6865
Detail $\surd$ 31.63 0.0075 0.3254 0.3398 98.28 0.0324 0.4027 0.8659 136.95 0.0613 0.3695 0.7938 60.40 0.0243 0.3659 0.6665